five years ago a women age is equal to the square of the age of her child,after 15 years her age is twice the age of her child.what is the present age of women?
Answer by jorel1380(3719)

Let c be the age of the child now. Five years ago, he was c-5, and his mother was (c-5)². Fifteen years from now:
(c-5)²+20=2(c+15)
c²-10c+45=2c+30
c²-12c+15=0
This yields a positive root of 10.582575695 as the age of the child now. and (c-5)²+5=36.1651513 as the age of the mother now. ☺☺☺☺